OpenStudy (anonymous):

If an organization lost 1.10 if its members . If 198 is left, how many members were there before the reduction?

OpenStudy (anonymous):

Do you mean 1 tenth? Or 1 and 1 tenth?

OpenStudy (anonymous):

1 tenth

OpenStudy (anonymous):

suppose before reduction there were total x members

OpenStudy (anonymous):

then organization lost x/10 memebers

OpenStudy (anonymous):

then u have to fill in the blank x- ...... = 198

OpenStudy (anonymous):

198/9=22 22*10=220 I believe the final answer is 220.

OpenStudy (anonymous):

x-22=198

OpenStudy (anonymous):

Yes..@Mello is correct

OpenStudy (anonymous):

Ok, that answer is correct mello. I understand now

OpenStudy (anonymous):

Try this.... A compamy will need $25,000 in7 years for a new addition. To met the goal, the company deposits money in an account today that pay 3% annula interest compund quarterly. Find the amount that should be invested to total %25,000 in 7 years?